OpenStudy (anonymous):

how do you get the radius of 5 with the points (1/rad.5, -2/rad5)

OpenStudy (tkhunny):

\(\left(\dfrac{1}{\sqrt{5}}\right)^{2} + \left(\dfrac{-2}{\sqrt{5}}\right)^{2} = \dfrac{1}{5} + \dfrac{4}{5} = 1\) Answer to your question is, "you don't". Why told you that you could?

OpenStudy (anonymous):

the question is in standard position and the terminal side of theta intersects the unit circle at the point (1/rad5,-2/rad5) find sin cos and tan

OpenStudy (tkhunny):

What Quadrant is it in?

OpenStudy (anonymous):

it doesn't say

OpenStudy (tkhunny):

Yes it does. Where does it intersect the unit circle? That tells you.

OpenStudy (anonymous):

its in standard position and it only gives you the points i listed above

OpenStudy (tkhunny):

And that tells you the quadrant. Come on. You have to get this piece. x > 0 and y < 0. What Quadrant is it in?